Systems Staff Meeting Notes 03/11/02

Systems Staff Meeting Notes
March 11, 2002

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart
  1. Take home test – group discussion on the location and function of the servers on the Smatherslib domain. In the event of a power outage or missing personnel, everyone needs to be aware of the servers and their connection with particular areas and problems.
  2. Geoffrey had a good time on his vacation.
  3. Ying had a girl – 7lbs, named Celina.
  4. Cynthia saw the trailer for the new Star Wars movie. It was good.
  5. The Exchange server died this morning – disk space, RAID failure. Will has installed one of the new replacement drives. Had to move the Office 2K share around to a different volume.
  6. Office Scan – was in mobile mode on Winston’s machine – they updated the engine over the weekend.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Dells not allocated yet. Jami Beserock needs a new computer & she is on Marvin’s list. Marvin already has one to give her.
  2. The memory has arrived.
  3. SAN order should (really) be going out this week.
  4. Microtek 9600XL scanner for the LAD project – was just received this morning. Means the LAD project should be able to get underway this week.
  5. Epson scanners – working fine after downloading the correct drivers. Bobbie replaced an HP scanner for Elena Aissing on Friday afternoon. There are no more left in the vault. Geoffrey notes that Grace still has a non-functional HP.
  1. Public W2K rollout – new rules to comply with, per DDD memo on Wednesday afternoon. Everything has to be W2K or it’s off the network. There was some confusion on Course Reserves – Bobbie is working with Tom Caswell – he is hesitant, is waiting on a scanner. Bill wants a list of replacement scanners by this afternoon.
  2. Xerox Print Release stations – are crippled. Xerox controls, they need to get with the program. No one has heard from the Xerox representative.
  3. Kiosk mode – 3270 sessions work signing in, but it requires more work. It comes up in Hummingbird - open profile - and won’t restart automatically at exit. It is useable in its current state.
  1. Trend Virus Updates – Geoffrey says the public machines have to shut down each time in order to update. Current scan engine should be 6.11. The machines are not picking up the auto update. Need to check all – not just public workstations.
  2. Power Outage – caused the loss of the NT2 HD.
  3. Network Browse – better, but not perfect. PDC is registering all, but not broadcasting back out. May want to check to make sure that the port is not blocked.
  4. Real Users (staff) do not belong in Power Users. Outlook will not let you sign-on the first time unless you are in Power Users. We have been doing that for staff to be able to run Outlook, print, etc. Outlook dies without Power User authority. We do not want it to be set up like this unless there is no further way to make it operational. Need to change the installation program to install now, rather than on first use.
Software Projects Status
  1. DLC post Ying – projects currently on hold. Need an update as to when Ying is expected to return. She’ll probably be taking six weeks off.
  2. No applet for kiosks – Hummingbird will do what we want. Hummingbird needs to be updated on every public machine.
  3. NT4 – Winston replaced the operating system, but could not upgrade SQL to the Enterprise version. Lack of terminal services prevented it from loading on the machine. Terminal service can be made available but not running & it will probably upgrade. Winston will try again either today or tomorrow – after 2AM.
  4. Event log – error message on NT4 – lack of LMHost file.
  5. FilmLog – Hilton has received several requests to add on to Phase 2 from Gus (Phase 2, Part 2). Changes should be finished today. All project-related time is being tracked.
  6. Personnel – details of OPS changes are being worked out.
  7. Net Agent – has been upgraded to 5.2. We’ve been getting requests for transcripts, now it is automatically sent out with this version.
DP Coordinator
  1. Students have returned from Spring Break.
  2. Students should be paging people when they are not at their desk but logged in to the office.
  3. PAULA – remember to give a detailed account – students do not know acronyms.
  4. Bill’s pager numbers have been updated. We keep everyone’s numbers on NT2R. Make sure the list you have is current.

Back to the Systems Homepage

Last updated March 25, 2002
by Debra Harris